OpenStudy (meilendurcer):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (meilendurcer):

@Salmon please help!

OpenStudy (anonymous):

Just multiply equation one by 2 and then substract it from equation two You can easily get your answer...

OpenStudy (meilendurcer):

I am not good at this math but i think it's either A or D

OpenStudy (anonymous):

Its D... x+3y=-1 .......(1) multipling by 2 2x+6y=-2 ........(2) 2x+2y=6 ..........(3) (2)-(3) 4y=-8 y=-2 Substitute y=-2 in equation (1) x-6=-1 x=5 so answer is D (5,-2)

OpenStudy (meilendurcer):

Yay!! also can you check my other ones? i have the questions and answers and want to make sure theyre right.

OpenStudy (anonymous):

Sure....